-4 square root(x+9)=20
square root of (x+9)= -5
we got this by dividing -4 on both sides,
now squaring on both sides
x+9 = 25
so
x+9-9=25-9
x=16.

Answer by sravancv(7)

-4 square root(x+9)=20
square root of (x+9)= -5
we got this by dividing -4 on both sides,
now squaring on both sides
x+9 = 25
so
x+9-9=25-9
x=16.